Two point charge placed at a distance 'r' in air exert a force 'F'. The distance at which they exert same force when placed in certain medium (dielectric constant K) is
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Since the force in both cases is same. So,
(1/4*pi* epsilon) ( q1 *q2/r²) = (1/4*pi* epsilon) (q1*q2/k *r'²)
1/r² = 1/k * r'²
r' = r/sqrt(k)
Regards
Arun (askIITians forum expert)